OpenStudy (anonymous):

Please guide me c: Find the area enclosed by the curves \(\sf y= \sqrt[3]{x}, y =1, y=2\ and x=0.\)

OpenStudy (anonymous):

So first, i'm gonna draw a graph

OpenStudy (anonymous):

thenn ?

OpenStudy (mathmale):

Nice drawing. Great way to start. Going to divide this area into strips? If so, will they be vertical or horizontal, and why?

OpenStudy (xapproachesinfinity):

use horizontally simple better for taking the integral

OpenStudy (mathmale):

xapproaches: why not let Data_LG2 answer that question, which was meant for him/her?

OpenStudy (xapproachesinfinity):

just suggest! of course he/she use whatever seen fit :)

OpenStudy (anonymous):

it will be horizontal? sorry, i had to refresh my browser.it is sticking.. brb

OpenStudy (xapproachesinfinity):

no it can be anything you want it was just a suggestion try to use your own approach, wht you see better

OpenStudy (solomonzelman):

\(\int\) f(x) - g(x) dx This is what I would use for x=[1,8] And for x=[0,1] it is a rectange.

OpenStudy (mathmale):

Good choice. Having made that choice, how would you represent the length of a given horizontal strip algebraically?

OpenStudy (anonymous):

now i'm back. so if I'll use the horizontal, i have to do something like this: \(\sf f(x) = \sqrt[3]{x}\) --> \(\sf x=f^{-1}(y)=y^3\) to make the integral easier to evaluate.... just thinking about using it, not really sure

OpenStudy (anonymous):

my interval will be [1,2]

OpenStudy (xapproachesinfinity):

yep :)

OpenStudy (xapproachesinfinity):

you are going very well

OpenStudy (mathmale):

Your strips are horiz., so you figure their lengths as the difference of two x values. Consider how you'd represent those two x values. Indicate which is the larger of the two.

OpenStudy (mathmale):

Also consider: are you integrating with respect to x or to y?

OpenStudy (anonymous):

with respect to y

OpenStudy (mathmale):

Agreed. Because that's the case, your limits of integration, 1 to 2, are fine.

OpenStudy (xapproachesinfinity):



OpenStudy (mathmale):

So, please write an expression for the AREA of a typical horiz. strip. What is the thickness of this strip? What is the length?

OpenStudy (xapproachesinfinity):

 just visualizing the strips :)